Improper Integral (Integrals with Infinite Limits)


∫e^x dx from - ∞ to 1


1
Expert's answer
2021-05-07T10:23:17-0400

Given Improper integral is-


"I=\int_{-\infty}^1e^xdx"


So, This integral can be written as-

"I=lim_{p\to \infty}\int_-p^1e^xdx



 \\ =lim_{p\to \infty}e^x|_p^1



 \\ =lim_{p\to \infty} (e-e^{-p})\\"


"=e-\dfrac{1}{e^{\infty}} =e-\dfrac{1}{\infty}=e-0=e"